g(x)=x^2-10x-3
a. g(-1) b. g(x+2) c. g(-x)

Found 2 solutions by themasterofcircuits, Tatiana_Stebko:
Answer by themasterofcircuits(27)

g(x) = x^2-10x-3
a.)
g(-1) = (-1)^2-10(-1)-3 = 1+10-3=8
b.)
g(x+2) = (x+2)^2-10(x+2)-3 = x^2+4x+4-10x-20-3 = x^2-6x-19
c.)
g(-x) = (-x)^2-10(-x)-3 = x^2+10x-3

Answer by Tatiana_Stebko(1539)